### v2

# Pre-DPO: Improving Data Utilization in Direct Preference Optimization Using a Guiding Reference Model

Abstract
Direct Preference Optimization (DPO) simplifies reinforcement learning from human feedback (RLHF) for large language models (LLMs) by directly optimizing human preferences without an explicit reward model. We find that during DPO training, the reference model plays the role of a data weight adjuster. However, the common practice of initializing the policy and reference models identically in DPO can lead to inefficient data utilization and impose a performance ceiling. Meanwhile, the lack of a reference model in Simple Preference Optimization (SimPO) reduces training robustness and necessitates stricter conditions to prevent catastrophic forgetting. In this work, we propose Pre-DPO, a simple yet effective DPO-based training paradigm that enhances preference optimization performance by leveraging a guiding reference model. This reference model provides foresight into the optimal policy state achievable through the training preference data, serving as a guiding mechanism that adaptively assigns higher weights to samples more suitable for the model and lower weights to those less suitable. Extensive experiments on AlpacaEval 2.0 and Arena-Hard v0.1 benchmarks demonstrate that Pre-DPO consistently improves the performance of both DPO and SimPO, without relying on external models or additional data. [^1] Our code and data can be found at https://github.com/DtYXs/Pre-DPO.
1 Introduction
Reinforcement learning from human feedback (RLHF) has become a crucial paradigm in aligning large language models (LLMs) with human preferences and values (Christiano et al., 2017; Ouyang et al., 2022; Achiam et al., 2023). This approach involves training an explicit reward model to provide reward signals and performing reinforcement learning under a KL constraint by a reference model instantiated identically to the initial policy. Direct Preference Optimization (DPO) (Rafailov et al., 2023), as a representative of reference-based preference optimization methods (Rafailov et al., 2023; Ethayarajh et al., 2024; Azar et al., 2024; Gorbatovski et al., 2024), transforms the objective of RLHF into direct preference training of the language model, enabling offline preference learning under the constraint of the reference model.
Recently, it has been shown that a reference model is not necessary for achieving effective preference optimization. Simple Preference Optimization (SimPO) (Meng et al., 2024), as a representative of reference-free preference optimization methods (Xu et al., 2023; Xu et al., 2024; Hong et al., 2024; Meng et al., 2024), eliminates the need for a reference model and yields better performance and efficiency, though at the cost of an increased risk of catastrophic forgetting (Meng et al., 2024). Meanwhile, other studies have empirically demonstrated that DPO can benefit either from relaxing the constraints imposed by the reference model (Gorbatovski et al., 2024) or from stronger external reference models (Liu et al., 2024). However, decreasing reliance on a reference model imposes stricter practical requirements to ensure effective learning (Meng et al., 2024), and clear methodologies for obtaining an appropriate reference model are still lacking.
Despite the empirical efforts, the role of the reference model in DPO and its impact on training dynamics remain insufficiently explored. We find that during DPO training, the reference model plays the role of a data weight adjuster (see Section 4.1). It adaptively tends to assign higher weights to data aligned with itself while reducing weights for conflicting data. However, due to the common practice in DPO of initializing the policy and reference models identically (Rafailov et al., 2023), as training progresses, the reference model increasingly constrains the policy model by penalizing deviations, potentially introducing a performance ceiling. Moreover, identical initialization of policy and reference models results in the nearly uniform weighting of training examples during the early training stages, which is in contrast to prior studies showing that assigning non-uniform weights to training data can lead to improved learning and performance (Lin et al., 2017; Ren et al., 2018; Shu et al., 2019).
In light of the limitations of conventional reference models, we hypothesize that an effective reference model for DPO should (1) differ from the initial policy model and (2) provide insights into potential directions for policy improvement based on the preference data. We define this type of reference model as a guiding reference model, which can better support learning by transforming the role of the reference model from a static constraint into a guide with foresight (see Section 4.2). Building on this insight, we propose Pre-DPO, a simple yet effective training paradigm that enhances data utilization and improves the performance of existing preference optimization methods without relying on external models or additional data. Pre-DPO first optimizes the initial policy using a standard preference optimization method (e.g., DPO or SimPO). The resulting optimized policy is then employed as the guiding reference model. Finally, the initial policy is re-optimized using DPO with this guiding reference model, yielding a better-optimized policy through more effective data reweighting. The guiding reference model in Pre-DPO essentially serves as an adaptive guiding mechanism that dynamically assigns higher weights to samples more suitable for the model and lower weights to those less suitable. Specifically, these suitable cases typically correspond to examples that are easier to learn, allowing the model to leverage data that aligns well with its learning trajectory efficiently.
Figure 1: Pre-DPO further enhances the performance of DPO and SimPO by leveraging a guiding reference model, consistently across different models and scales.
Our extensive experiments results conducted on the Llama3.2 (Grattafiori et al., 2024) and Qwen2.5 (Yang et al., 2024) model series demonstrate that Pre-DPO can consistently improve the performance of both DPO and SimPO on the AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard v0.1 (Li et al., 2024) benchmarks, achieving average improvements of 2.5 points in the AlpacaEval 2 length-controlled win rate (LC) and 2.6 points in the Arena-Hard v0.1 win rate (WR). By introducing the guiding reference model, Pre-DPO can further improve the performance of existing well-tuned preference optimization methods (as illustrated in Figure 1), effectively overcoming the performance ceiling caused by inefficient data utilization under traditional reference model settings. Notably, Pre-DPO does not rely on external models or additional data, making it highly flexible and easy to deploy.
2 Related work
RLHF is currently an effective technique for aligning LLMs with human values and preferences (Ouyang et al., 2022; Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024; Shen et al., 2025). Typically, pretrained LLMs first undergo SFT to acquire the ability to follow human instructions. Subsequently, preference data is annotated by human or AI annotators to train a reward model. Finally, policy optimization is performed using reinforcement learning methods, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), Group Relative Policy Optimization (GRPO) (Shao et al., 2024), REINFORCE++ (Hu, 2025), etc. Although RLHF has demonstrated superior performance, the entire pipeline is complex and challenging.
In contrast, offline preference optimization eliminates the need for an explicit reward model and avoids the complex online reinforcement learning optimization process. A well-optimized model obtained via preference optimization can also serve as a strong initial policy for subsequent reinforcement learning optimization (Yang et al., 2024). Depending on whether or not a reference model is required, preference optimization can be classified into two categories: reference-based preference optimization methods (Rafailov et al., 2023; Ethayarajh et al., 2024; Azar et al., 2024; Gorbatovski et al., 2024) and reference-free preference optimization methods (Xu et al., 2023; Xu et al., 2024; Hong et al., 2024; Meng et al., 2024). DPO (Rafailov et al., 2023), as a representative of reference-based preference optimization methods, directly trains on preference data and implicitly optimizes the same objective as existing reinforcement learning algorithms. SimPO(Meng et al., 2024), as a representative of reference-free preference optimization methods, removes the need for a reference model and can achieve better results than DPO at the cost of lower training robustness (Meng et al., 2024). In this paper, we experimentally demonstrate that DPO can also benefit from reference-free preference optimization methods through guiding reference models.
For the reference models in DPO, prior work (Liu et al., 2024) empirically suggested that DPO can benefit from a stronger and more suitable reference model in certain cases. However, it mainly focuses on stronger external reference models and does not provide a theoretical explanation for why they can be beneficial. Gorbatovski et al. (2024) proposed a dynamic update strategy to reset the reference model based on the current policy, which could weaken the regularization effect of the reference model and tends to assign more identical weights to the data samples with more frequent updates. In this work, we introduce the concept of a guiding reference model, analyze its impact on maximizing data utilization in DPO through data reweighting, and propose an effective methodology for leveraging a guiding reference model in DPO.
3 Preliminaries
Reinforcement learning from human feedback (RLHF).
Given a text prompt $x$, the RLHF training stage aims to increase the probability that an LLM generates a response $y$ that is better aligned with human values. Specifically, the objective is to maximize the expected reward $r(x,y)$ while controlling the deviation between the policy probability distribution $\pi_{\theta}(y\mid x)$ and the reference probability distribution $\pi_{\text{ref}}(y\mid x)$. The optimization objective is formulated as follows:
$\max_{\pi_{\theta}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{\pi_{\theta}}}\left[r(x,y)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(y\mid x)\parallel\pi_{\text{ref}}(y\mid x)\right],$ (1)
where $\mathcal{D}_{\pi_{\theta}}$ denotes the distribution of the policy model. The KL divergence term constrains the deviation of the policy $\pi_{\theta}$ from the reference model, and $\beta$ controls the strength of this constraint.
Direct Preference Optimization (DPO).
DPO (Rafailov et al., 2023) is a widely used reference-based preference optimization method that eliminates the need for explicit reward signals and has become a component in the post-training pipeline of many popular open-source LLMs (Bi et al., 2024; Jiang et al., 2024; Yang et al., 2024; Xu et al., 2025). It reformulates Eq. 1 into a direct optimization process on a preference dataset $\mathcal{D}=\{(x_{i},y^{+}_{i},y^{-}_{i})\}_{i=1}^{|\mathcal{D}|}$, where $x$ is the prompt, $y_{i}^{+}$ is the preferred response, and $y_{i}^{-}$ is the less-preferred response. The objective of DPO is as follows:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\Big{(}\beta\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}-\beta\log\frac{\pi_{\theta}(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\Big{)}\right],$ (2)
where $\sigma(\cdot)$ is the sigmoid function, and $\beta$ is a hyperparameter that controls the strength of the reference model’s constraint.
A large $\beta$ imposes a strong constraint, which can limit the model’s performance improvement, while a small $\beta$ may result in insufficient constraints, leading to model degradation (Liu et al., 2024). In practical LLM training, $\pi_{\theta}$ and $\pi_{\text{ref}}$ are typically initialized as the supervised fine-tuning (SFT) model. $\pi_{\text{ref}}$ remains fixed during training.
Simple Preference Optimization (SimPO).
SimPO (Meng et al., 2024) is a reference-free preference optimization method that removes the reference model in DPO while introducing length normalization and target reward margin. Its loss function is defined as:
$\mathcal{L}_{\text{SimPO}}(\pi_{\theta})=-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\log\sigma\left(\frac{\beta}{|y^{+}|}\log\pi_{\theta}(y^{+}\mid x)-\frac{\beta}{|y^{-}|}\log\pi_{\theta}(y^{-}\mid x)-\gamma\right)\right],$ (3)
where $|y^{+}|$ and $|y^{-}|$ denote the lengths, $\beta$ is a hyperparameter constant, and $\gamma$ is the target reward margin. SimPO has the potential to surpass DPO in performance but suffers from reduced robustness due to the lack of reference constraints (Meng et al., 2024).
4 Method
Figure 2: An overview of Pre-DPO. DPO constrains training using the initial policy model as the reference, and SimPO is reference-free. Pre-DPO first obtains an optimized policy model via DPO or SimPO, resets it as a guiding reference model, and then re-optimizes the initial policy using DPO. This process improves data utilization and leads to a better optimized policy model.
In this section, we first present the motivation behind Pre-DPO, focusing on the limitations of the reference setting in DPO and the specific challenges Pre-DPO addresses. Then, we provide a detailed explanation of Pre-DPO, outlining how it adaptively reweights the training examples and describing its overall process (as shown in Figure 2).
4.1 The reference model in DPO is a data weight adjuster
From the loss function of DPO (Eq. 2), we can derive the gradient with respect to the parameters $\theta$:
$\nabla_{\theta}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\beta\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\left[\lambda\cdot\nabla_{\theta}\log\frac{\pi_{\theta}(y^{+}\mid x)}{\pi_{\theta}(y^{-}\mid x)}\right],$ (4)
where $\lambda$ is defined as:
$\lambda=\sigma\left(\beta\log\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}-\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\theta}(y^{-}|x)}\right).$ (5)
From the perspective of example reweighting, DPO learns from preference pairs with weights $\lambda$, where the reference model $\pi_{\text{ref}}$ controls the training process by adjusting $\lambda$.
When the policy $\pi_{\theta}$ and the reference $\pi_{\text{ref}}$ are initialized from an identical SFT model, $\lambda$ starts around the constant 0.5 in the early training period due to $\sigma(0)=0.5$. However, a constant $\lambda$ can lead to degeneracy (Rafailov et al., 2023), and more importantly, previous research in the field of example reweighting has shown that assigning appropriate and varying weights to training samples can improve model performances and data efficiency (Lin et al., 2017; Ren et al., 2018; Shu et al., 2019).
On the other hand, as training progresses, the reference continuously constrains the policy’s deviation by adjusting the value of $\lambda$. Specifically, when $\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}$ is large, it encourages a larger value of $\lambda$, promoting learning from the corresponding preference pair. A small ratio typically results in a reduced value of $\lambda$, which can reduce the model’s learning from that sample. Therefore, a suboptimal configured reference model can lead to suboptimal weighting of training samples.
4.2 Pre-DPO: improving data utilization in DPO using a guiding reference model
One straightforward solution is to employ a reference model that differs from the initial policy and provides foresight into promising directions for policy improvement based on the preference data $\mathcal{D}$, enabling more effective data reweighting and guidance during training.
Notably, a model already undergone preference optimization contains information about the entire training process. More importantly, it reflects the outcomes that the initial policy can achieve through the available preference data. Specifically, when the reference model in DPO is set to a guiding reference model $\pi_{\text{guide}}$, the weight $\lambda$ becomes:
$\lambda=\sigma\left(\beta\log\frac{\pi_{\text{guide}}(y^{+}|x)}{\pi_{\text{guide}}(y^{-}|x)}-\beta\log\frac{\pi_{\theta}(y^{+}|x)}{\pi_{\theta}(y^{-}|x)}\right).$ (6)
The foresight of the guiding reference model is embodied in its ability to dynamically modulate $\lambda$, assigning higher weights to samples that the policy model is capable of learning well and down-weighting those that are difficult to learn or even conflict with the model’s learning trajectory. This behavior naturally satisfies the conditions suggested in the concurrent work (Gao et al., 2025), where avoiding overly difficult examples benefits alignment. Compared to the reference model used in standard DPO, which applies a fixed constraint with no foresight, the guiding reference model offers adaptive and data-dependent reweighting, leading to more efficient and targeted policy improvement.
Hence, we propose a simple yet effective paradigm for obtaining a suitable reference without needing external models. Specifically, we reset the optimized policy model as the reference for the next training iteration. Since the reference model retains all the information from prior policy training, its role shifts from a constraint to that of a guide, which we refer to as the guiding reference model. Employing this guiding reference model in DPO adaptively assigns higher weights to training data that aligns with it while reducing the weights of conflicting samples.
Let $\pi_{\theta}$ denote the policy model to be optimized, $\pi_{\text{ref}}$ represent the reference model (which can also be set as None), and $\mathcal{M}(\pi_{\theta};\pi_{\text{ref}})$ indicate the preference optimization method. The procedure of Pre-DPO (illustrated in Figure 2) is described in detail below:
Step1: Instantiate the initial reference model.
If $\mathcal{M}$ is reference-based, set the reference $\pi_{\text{ref}}$ to $\pi_{\text{SFT}}$. Otherwise, for reference-free methods, $\pi_{\text{ref}}$ is set to None:
$\pi_{\text{ref}}=\begin{cases}\pi_{\text{SFT}}&\text{for reference-based $\mathcal{M}$},\\ \texttt{None}&\text{for reference-free $\mathcal{M}$}.\end{cases}$ (7)
Step2: The first round of preference optimization.
Perform preference optimization $\mathcal{M}$ on $\pi_{\text{SFT}}$ with the preference dataset $\mathcal{D}$:
$\pi_{\mathcal{M}}=\mathcal{M}(\pi_{\text{SFT}};\pi_{\text{ref}}).$ (8)
Step3: Set the guiding reference model.
After the first round of optimization, reset $\pi_{\text{ref}}$ to the optimized model $\pi_{\mathcal{M}}$ obtained from the previous round. This optimized model now serves as the guiding reference model $\pi_{\text{guide}}$:
$\pi_{\text{guide}}=\pi_{\mathcal{M}}.$ (9)
Step4: Preference optimization with the guiding reference model.
Apply DPO to $\pi_{\text{SFT}}$ using the guiding reference $\pi_{\mathcal{\text{guide}}}$ on the same dataset $\mathcal{D}$ to obtain the better optimized model $\pi_{\text{Pre-DPO{}}}$:
$\pi_{\text{Pre-DPO{}}}=\mathcal{M}_{\text{DPO}}(\pi_{\text{SFT}};\pi_{\text{guide}}).$ (10)
5 Experiments
We empirically evaluate the effectiveness of Pre-DPO in enhancing existing preference optimization methods through a guiding reference model. To ensure a comprehensive and fair assessment, we conduct experiments on the Llama3.2 (Grattafiori et al., 2024) and Qwen2.5 (Yang et al., 2024) model series, including both Base and Instruct versions. We evaluate performance on two widely-used preference optimization benchmarks: AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard v1.0 (Li et al., 2024). Given the sensitivity of preference optimization to hyperparameters (Meng et al., 2024), we follow SimPO’s experimental methodology and conduct an extensive hyperparameter search to ensure reliable results. All experiments are conducted based on the LlamaFactory (Zheng et al., 2024) repository, and all models and datasets used are publicly available.
5.1 Experimental setup
Models.
In our experiments, we primarily consider two widely recognized series of open-source models, Llama3.2-3B and Qwen2.5-7B, including both Base and Instruct versions. The diversity in model types and scales enables a more comprehensive evaluation of the method’s effectiveness.
Datasets.
For Base models, we obtain their corresponding SFT versions by training on the UltraChat-200k (Ding et al., 2023) dataset [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k. Subsequently, the binarized version of the UltraFeedback (Cui et al., 2024) dataset [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized is utilized for preference optimization.
For Instruct models, we use them directly as SFT models. In the preference optimization stage, we construct two more on-policy preference datasets for Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct, respectively. Specifically, we use prompts from the UltraFeedback dataset to sample responses from the model, generating 6 responses per prompt with a temperature of 0.8 and top-p of 0.95. Subsequently, following previous work(Meng et al., 2024), we utilize the ArmoRM-Llama3-8B-v0.1 (Wang et al., 2024) reward model to score each response and select the highest-scoring and lowest-scoring responses to form preference pairs $(x,y^{+},y^{-})$.
Implementation details.
For the SFT stage in Base models, we train for 3 epochs using a batch size of 32, a maximum sequence length of 4096, a learning rate of $2\times 10^{-6}$, and a cosine learning rate schedule with a 6% warmup ratio.
We obtain the guiding reference models for Pre-DPO based on two popular preference optimization methods: DPO (Rafailov et al., 2023), representing reference-based methods, and SimPO (Meng et al., 2024), representing reference-free methods. All preference optimization experiments use a batch size of 128, a maximum sequence length of 4096, and a cosine learning rate schedule with a 6% warmup ratio, training for 1 epoch. Given the importance of hyperparameter tuning in offline preference optimization (Meng et al., 2024), we perform extensive hyperparameter searches for all preference optimization experiments to ensure fairness. Specifically, for the key hyperparameters, including the learning rate, $\beta$ (for DPO and SimPO), and $\gamma$ (for SimPO), a two-stage tuning strategy is employed. We first fix the learning rate and search for the optimal $\beta$ or $\gamma$. Then, with the best $\beta$ or $\gamma$ fixed, we search for the optimal learning rate. More details of hyperparameter tuning and the best hyperparameter setting in our experiments can be found in Appendix A.1.
Evaluation benchmarks.
We evaluate methods primarily on two open-source instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard v0.1 (Li et al., 2024), which are widely adopted in the community for evaluating the instruction-following capabilities of LLMs (Meng et al., 2024). We report the raw win rate (WR) and the length-controlled win rate (LC) on AlpacaEval 2, and the win rate (WR) on Arena-Hard v0.1, with GPT-4-Turbo serving as the judge model. More evaluation details can be found in Appendix A.2.
5.2 Main results
Method Ref. Llama3.2-3B-Base Llama3.2-3B-Instruct
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC (%) WR (%) Len. WR (%) LC (%) WR (%) Len. WR (%)
SFT - 6.1 4.0 1012 2.1 19.0 18.9 1956 18.5
DPO SFT 10.5 12.0 2042 10.6 36.3 36.9 2026 30.6
SimPO - 13.1 13.1 1950 11.7 33.8 29.9 1797 28.1
Pre-DPO DPO 12.5 13.9 2061 11.9 39.3 40.9 2095 34.7
Pre-DPO SimPO 18.1 18.4 2020 14.0 35.0 32.3 1846 30.0
Method Ref. Qwen2.5-7B-Base Qwen2.5-7B-Instruct
AlpacaEval 2 Arena-Hard AlpacaEval 2 Arena-Hard
LC (%) WR (%) Len. WR (%) LC (%) WR (%) Len. WR (%)
SFT - 18.6 6.9 892 9.4 31.2 31.0 2020 55.9
DPO SFT 24.8 22.2 1778 33.1 52.2 56.8 2270 62.9
SimPO - 34.7 31.9 1836 38.1 51.7 52.4 2119 62.4
Pre-DPO DPO 27.4 24.5 1790 32.6 53.3 59.4 2322 68.8
Pre-DPO SimPO 37.2 32.6 1758 41.6 54.6 55.5 2121 64.5
Table 1: Performance of Pre-DPO under four different model settings on AlpacaEval 2 and Arena-Hard v0.1. LC and WR denote the length-controlled and raw win rate, respectively. Ref. denotes the reference model and Len. denotes the average response length. The SFT models for the Base settings are trained on the UltraChat-200k dataset, while the Instruct models are used as the SFT models directly for the Instruct settings. The guiding reference models are obtained from DPO and SimPO.
Pre-DPO further improves DPO and SimPO by leveraging guiding reference models.
In Table 1, we report the performance on the AlpacaEval 2 and Arena-Hard v0.1 across the Llama3.2-3B-Base, Llama3.2-3B-Instruct, Qwen2.5-7B-Base, Qwen2.5-7B-Instruct settings. Compared with baselines, Pre-DPO achieves better performance on the AlpacaEval 2 LC and WR benchmarks, yielding average improvements of 2.5 and 2.8 points, respectively. On the Arena-Hard v0.1 benchmark, Pre-DPO also consistently demonstrates improvements across most settings. For instance, on Qwen2.5-7B-Instruct, Pre-DPO achieves an improved WR of 68.8 compared to 62.9 of the DPO baseline. These results indicate that Pre-DPO is effective in further improving both reference-based and reference-free methods by leveraging guiding reference models.
Pre-DPO improves performance without significantly increasing the response length.
Although Pre-DPO continuously improves performance, we observe that it does not significantly increase the response length compared to the baselines. Notably, in the Qwen2.5-7B-Base setting with SimPO as the guiding reference model, Pre-DPO even achieves the best performance with the shortest response length.
Additional observations.
Interestingly, we observe that SimPO shows a significant advantage over DPO in the Base settings, while DPO slightly outperforms SimPO in the Instruct settings. A possible reason is that the preference datasets constructed via on-policy sampling exhibit smaller length differences between positive and negative responses (see Appendix B.1 for details), thus diminishing the advantage of SimPO’s length normalization.
5.3 Ablations
The guiding reference model plays a critical role in the improvement of Pre-DPO.
Policy Model Method Ref. Epochs AlpacaEval 2
LC (%) WR (%) Len.
Llama3.2-3B-Base-SFT DPO SFT 1 10.5 12.0 2042
DPO SFT 2 11.0 12.0 1976
Pre-DPO $\text{DPO}_{1}$ 1 12.5 13.9 2061
Llama3.2-3B-Instruct DPO SFT 1 36.3 36.9 2026
DPO SFT 2 35.2 37.1 2113
Pre-DPO $\text{DPO}_{1}$ 1 39.3 40.9 2095
Table 2: Ablation studies of the guiding reference model. DPO trained for 2 epochs has the same computational cost as Pre-DPO. $\text{DPO}_{1}$ denote the guiding reference model trained with DPO for 1 epoch.
Although Pre-DPO consistently improves the performance of DPO and SimPO, it introduces additional computational cost due to the need to obtain a guiding reference model. To investigate whether the performance gain is simply due to increased training, we compare it with a baseline that trains DPO for 2 epochs with the original reference setting. As shown in Table 2, under the same computational budget, Pre-DPO with a guiding reference model consistently achieves the best LC and WR on AlpacaEval 2, benefiting from the better data utilization enabled by the guiding reference model and avoiding the excessive constraints imposed by traditional reference model setups.
Guiding reference models perform better on previously encountered preference data.
Method Ref. Training Set AlpacaEval 2
LC (%) WR (%) Len.
DPO SFT A 11.6 10.5 1702
Pre-DPO $\text{DPO}_{\text{A}}$ A 13.9 13.2 1819
Pre-DPO $\text{DPO}_{\text{B}}$ A 12.9 12.5 1791
DPO SFT B 13.9 12.9 1765
Pre-DPO $\text{DPO}_{\text{B}}$ B 14.1 13.4 1792
Pre-DPO $\text{DPO}_{\text{A}}$ B 13.8 12.7 1708
Table 3: Ablation studies of Pre-DPO under different guiding reference models and preference datasets. The SFT model is obtained by training Qwen2.5-3B-Base on the UltraChat-200k dataset. The preference training sets A and B are created by randomly splitting the UltraFeedback dataset into two equal parts. $\text{DPO}_{\text{A}}$ and $\text{DPO}_{\text{B}}$ denote the guiding reference models trained on training sets A and B, respectively.
To investigate the optimal relationship between the guiding reference model and the preference training data, we divide the preference dataset into two equal parts, A and B. We then perform DPO on each part separately to obtain two distinct guiding reference models. For each subset of preference data, we apply Pre-DPO using both guiding reference models for comparison. The results, as shown in Table 3, indicate that a guiding reference model trained on the same subset of preference data yields better performance on AlpacaEval 2. This suggests that guiding reference models perform better when guiding policies on data they have previously encountered, as they possess more accurate information about potential policy optimization directions and enable more effective data weighting.
5.4 The $\lambda$ distribution after training of Pre-DPO and vanilla DPO
(a) Llama3.2-3B-Base
(b) Llama3.2-3B-Instruct
Figure 3: Comparative differences in $\lambda$ (when $\beta$ is set to 0.01) distributions between Pre-DPO and DPO across Llama3.2-3B settings. Each subplot presents a violin plot illustrating the distribution of $\lambda$ values on the training preference dataset. The black horizontal line indicates the mean, while the dark gray line represents the median. The width of the violin at each point represents the data density, where thicker sections correspond to regions with higher data concentration.
Based on Eq. 5, Pre-DPO demonstrates a distinct advantage through its adaptive data reweighting mechanism, which dynamically assigns higher weights to preference data that align with the learned outcomes of the guiding reference model while diminishing the influence (assign lower weights) of data that may introduce contradictory learning signals. Figure 3 illustrates the differences in $\lambda$ distributions on the training preference dataset between Pre-DPO and DPO on Llama3.2-3B settings. For clarity, we report the statistics of the static $\lambda$ values computed on the training dataset after training completion. Pre-DPO produces higher mean and median $\lambda$ values compared to DPO, indicating that the guiding reference model effectively guides the policy model toward better data utilization. More comparisons between Pre-DPO and the baselines are provided in Appendix B.2.
6 Conclusion
We proposed Pre-DPO, a simple yet effective DPO-based preference optimization paradigm that enhances data utilization and improves performance by leveraging a guiding reference model. Unlike traditional DPO, which uses a reference identical to the initial policy, Pre-DPO reuses a previously optimized guiding reference model to re-optimize the initial policy model. This shifts the reference model’s role from constraint to guidance, enabling better data reweighting. Extensive experiments demonstrate that Pre-DPO consistently outperforms both DPO and SimPO across various models and scales, without relying on external models or additional data. We hope this work can inspire more exploration and discussion on the role and improvement of reference models in LLM alignment.
Appendix A More experimental details
A.1 Training hyperparameters
Hyperparameter tuning.
Considering that hyperparameter tuning is crucial for preference optimization, we adopt a two-stage hyperparameter search method to ensure both fairness and efficiency. Specifically, in the first stage, we fix the learning rate at $6\times 10^{-7}$ and individually search for the optimal $\beta$ in $[0.005,0.01,0.05,0.1,0.2,0.5,1.0]$ for DPO, and search for $\beta$ in $[2.0,2.5]$ and $\gamma$ in $[0.3,0.5,1.0,1.2,1.4,1.6]$ for SimPO. In the second stage, we select the two best hyperparameter settings found in the first stage and individually search for the optimal learning rate in $[3\times 10^{-7},5\times 10^{-7},8\times 10^{-7},1\times 10^{-6}]$. The search ranges for the hyperparameters are chosen with reference to prior work (Meng et al., 2024).
The optimal hyperparameter values.
Table 4 shows the optimal hyperparameter values in our main experiments. It can be observed that Pre-DPO generally requires a larger $\beta$ compared to DPO.
Policy Model Method Ref. $\boldsymbol{\beta}$ $\boldsymbol{\gamma}$ Learning rate
Llama3.2-3B-Base-SFT DPO SFT 0.005 - $1\times 10^{-6}$
Llama3.2-3B-Base-SFT SimPO SFT 2.5 1.2 $1\times 10^{-6}$
Llama3.2-3B-Base-SFT Pre-DPO DPO 0.05 - $1\times 10^{-6}$
Llama3.2-3B-Base-SFT Pre-DPO SimPO 0.05 - $6\times 10^{-7}$
Llama3.2-3B-Instruct DPO SFT 0.05 - $6\times 10^{-7}$
Llama3.2-3B-Instruct SimPO SFT 2.5 1.0 $1\times 10^{-6}$
Llama3.2-3B-Instruct Pre-DPO DPO 0.05 - $1\times 10^{-6}$
Llama3.2-3B-Instruct Pre-DPO SimPO 0.1 - $1\times 10^{-6}$
Qwen2.5-7B-Base-SFT DPO SFT 0.005 - $8\times 10^{-7}$
Qwen2.5-7B-Base-SFT SimPO SFT 2.5 1.4 $8\times 10^{-7}$
Qwen2.5-7B-Base-SFT Pre-DPO DPO 0.2 - $8\times 10^{-7}$
Qwen2.5-7B-Base-SFT Pre-DPO SimPO 0.2 - $1\times 10^{-6}$
Qwen2.5-7B-Instruct DPO SFT 0.01 - $5\times 10^{-7}$
Qwen2.5-7B-Instruct SimPO SFT 2.5 1.2 $1\times 10^{-6}$
Qwen2.5-7B-Instruct Pre-DPO DPO 0.05 - $1\times 10^{-6}$
Qwen2.5-7B-Instruct Pre-DPO SimPO 0.1 - $6\times 10^{-7}$
Table 4: Optimal hyperparameters for different experiments.
A.2 Evaluation details
Following previous work (Meng et al., 2024), we adopt a sampling-based decoding strategy for AlpacaEval 2 with a temperature of 0.7, a top-p of 0.9, and a greedy decoding strategy for Arena-Hard v0.1. We evaluate the AlpacaEval 2 and Arena-Hard v0.1 results using their official repositories, respectively.
Appendix B More analysis
B.1 Response length differences
We define the normalized length difference between positive and negative responses in preference datasets as follows:
Normalized Length $\text{Normalized Length Difference}=\frac{|\text{len}(y^{+})-\text{len}(y^{-})|}{\max\left(\text{len}(y^{+}),\text{len}(y^{-})\right)},$ (11)
where $\text{len}(y)$ denotes the number of tokens in response $y$.
We compute this metric on three datasets: the original UltraFeedback preference data used for Base models, and the resampled UltraFeedback-Llama3.2 and UltraFeedback-Qwen2.5 preference datasets used for Instruct models. The original UltraFeedback dataset exhibits an average normalized length difference of 0.465, while the resampled on-policy datasets, UltraFeedback-Llama3.2 and UltraFeedback-Qwen2.5, have smaller values of 0.296 and 0.224, respectively. This indicates that the resampled datasets contain less variation in response length, potentially imposing more stringent optimization conditions for SimPO.
B.2 The $\lambda$ distribution after training of Pre-DPO and vanilla DPO / SimPO
(a) Qwen2.5-7B-Base
(b) Qwen2.5-7B-Instruct
(c) Llama3.2-3B-Base
(d) Llama3.2-3B-Instruct
(e) Qwen2.5-7B-Base
(f) Qwen2.5-7B-Instruct
Figure 4: More violin plots showing the distribution of $\lambda$ values on the training preference dataset for Pre-DPO and DPO / SimPO across different settings (with $\beta=0.01$).
More differences in $\lambda$ distributions on the training preference dataset between Pre-DPO and DPO / SimPO are shown in Figure 4. Similar to the case where DPO is used as the guiding reference model, Pre-DPO also produces higher mean and median $\lambda$ values when using SimPO as the guiding reference.